A 0.70 kg mass oscillates according to the equation x=0.22sin(5.8t) where x is in meters and t is in seconds.

What is the Kinetic and Potential Energy when X=11cm
Answer #1

we know total energy = 1/2 * kA^2

k = m*w^2

x = Asin(wt) compare above equation with equation

w = 5.8 rad/s  

m = 0.70 kg

A = 0.22 m

total energy = 1/2 * 0.70 * (5.8)^2 * 0.22 = 0.57 J

at x = 11 cm = 0.11 m

potential energy = 1/2 * k * x^2

PE = 0.14 J

KE = total energy - PE = 0.43 J